# CharacterGen: Efficient 3D Character Generation from Single Images with Multi-View Pose Canonicalization

### v2

@@ Abstract
Figure 1: CharacterGen is an efficient 3D character generation framework. It takes a single input image and generates high-quality 3D character mesh in a canonical pose with consistent appearance, suitable for downstream rigging and animation workflows.
1 Introduction
The digital content industry’s rapid evolution has made the creation of high-quality 3D content a pivotal aspect across various domains, including film, video gaming, online streaming, and virtual reality (VR). Although manually modeled 3D content can attain exceptional quality, the significant time and labor investment required presents a substantial bottleneck. Addressing this, there has been a notable influx of exciting research [54, 69, 47, 46, 60, 9, 34, 20, 41, 3, 64, 59, 63, 29, 28, 32] focused on generating 3D models from single images. This approach substantially lowers the barrier to entry for novice users, democratizing access to 3D content creation and potentially revolutionizing the field.
3D character models often feature complex articulations, leading to frequent self-occlusion in 2D images that pose significant challenges in reconstruction, generation, and animation. Moreover, these characters may assume a range of body poses, including some that are rare and challenging to accurately interpret, leading to a diverse yet imbalanced data domain. These complexities hinder the effective generation, rigging, and animating of such models. As a result, general 3D generation techniques [45, 6, 62, 26, 58, 35] and single-view 3D reconstruction methods [56, 14, 59, 63, 64, 29, 3] often fall short in delivering optimal outcomes. Prior research [4, 16, 21, 17, 25, 66, 67, 49, 50, 76, 61] has explored the use of parametric models of 3D human bodies [33, 42, 2, 24] as 3D priors. However, these methods are predominantly tailored to realistic human proportions and relatively tight clothing, limiting their applicability. This constraint is especially noticeable in the context of stylized characters, known for their exaggerated body proportions and complex clothing designs, which challenge the adaptability and effectiveness of these approaches.
@@ 1 Introduction
Our 3D character generation approach has two tightly interconnected stages: initially, lifting a single image to multiple viewpoints while simultaneously canonicalizing the input pose; following this, we proceed to reconstruct a 3D character using this canonical pose. This method is supported by two key insights: firstly, it incorporates established principles and successful techniques from recent advancements in controllable image generation [74, 68]; secondly, it overcomes the challenges associated with sparse-view reconstruction for 3D characters. By focusing on the canonical pose, in which the geometric and texture structures are more clearly defined and self-occlusion is minimized, our approach simplifies the task of reconstructing both geometry and texture from limited views. The first stage involves a diffusion-based, image-conditioned multi-view generation model [31, 60, 52, 55], adept at capturing and translating both the global and local features character features from the input image to the canonical pose, which further facilitates the generation of consistent canonical pose images across multiple views. The second stage employs a transformer-based, generalizable sparse-view reconstruction model [14]. This model is key to generating a coarsely textured 3D character model from the images produced in the first stage. We further refine the model’s texture resolution through projective texture mapping and Poisson Blending [43], achieving a detailed final model. Furthermore, generating characters in a canonical pose also significantly benefits downstream applications, such as rigging and animation. Our whole generation process takes less than 1 mintue.
To train our pipeline, we have compiled a multi-pose, multi-view character dataset, focusing on anime characters due to their widespread availability online, notably on platforms such as VRoid Hub 111. We have amassed a collection of 13,746 characters and rendered these from various viewpoints across multiple body poses. This extensive collection has been organized into a dataset that we refer to as Anime3D.
@@ 2 Related Work
This section mainly discusses related works on diffusion-based 3D objects and avatar generation. Our CharacterGen also adopts transformer-based reconstruction models [14, 23] for efficient 3D character generation. Space limitations preclude discussion of 3D human reconstruction works [49, 50, 66, 67].
2.1 Diffusion-Based 3D Object Generation
@@
Recently, diffusion methods have shown a strong ability to guide 3D object generation tasks in the past year. Pioneering work of DreamFusion [45] and SJC [58] utilize score distillation sampling (SDS) to provide gradient guidance from pre-trained 2D diffusion models for text-to-3D generation tasks. Magic3D [26] and Fantasia3D [6] utilize an implicit tetrahedral field to support rendering with high resolution in the refinement stage. ProlificDreamer [62] proposes VSD to distill gradient scores from a LoRA network to better learn the distribution of 3D objects. Zero123 [30] presents a novel diffusion model to generate multi-view images that conform to the input image with given camera poses. Magic123 [46] combines both SDS and Zero123 guidance for generating 3D objects from image prompts, and adopts reconstruction loss to enhance front-view texture quality. MVDream [53] and ImageDream [60] utilize multi-view diffusion models to provide highly consistent guidance in 3D object generation process. SyncDreamer [31] utilizes 3D-aware attention modules to achieve synchronized multi-view image generation. Various other works [20, 71, 39, 18, 70, 41, 10] employ 3D data to train diffusion models for direct 3D object generation, which are fast, but struggle with output diversity.
2.2 3D Avatar Generation
@@
Using strong human-body priors such as SMPL [33] and SMPL-X [42], it is possible to generate high-quality human avatars based on the general 3D generation methods. EVA3D [13] combines a GAN backbone with a pose-guided sampling method to generate high-quality 3D human avatars. AvatarCLIP [12] first solves text-to-human generation by utilizing the pre-trained CLIP model to guide optimization of geometry and color networks. Dreamavatar [4] and AvatarCraft [19] utilize SMPL to initialize the implicit human geometry used in the diffusion-guided generation process. DreamHuman [21] adopts ImGHum [2] as body priors and proposes a focus rendering mechanism to better reconstruct the detailed geometry of avatars. DreamWaltz [16] utilizes ControlNet [74] to provide pose guidance to finetuning the animated representation. AvatarVerse [72] and AvatarStudio [73] both utilize a DensePose-guided ControlNet in the generation process to circumvent the multi-face “Janus” problem and to support part geometry optimization. TeCH [17] supports image-prompt avatar generation by training an additional DreamBooth model [48] on the input image for the SDS guidance model. TADA [25] directly distills 2D diffusion models to optimize the normals and displacements of an SMPL body mesh.
@@ 3.1 Anime3D Dataset
3.1.1 Data Acquisition
Existing large 3D object datasets like Objaverse [8] or OmniObject3D [65] do not contain enough 3D stylized characters for our training purposes. Inspired by PAniC3D [7], we first collected a large dataset of nearly 14,500 anime characters from the VRoidHub [57] and then removed non-humanoid data, to leave 13,746 character models.
Figure 2: An example character from our Anime3D dataset from four different camera views, demonstrates how we organize the image pairs during training to extend UNet’s ability to determine a canonical pose.
Figure 3: Pipeline for generating four views of consistent images, showing how our IDUNet extracts local pixel-level features to strengthen the multi-view UNet.
3.1.2 Data Processing
We need to render all the objects into image format to fine-tune 2D diffusion models. We utilize the three-vrm [44] framework to render the posed characters. We obtain A-pose characters and posed characters to generate canonical pose and random pose image pairs. Details of pose settings can be found in our supplementary material. We normalize the bounding boxes of the character model to $0.5,0.5]^{3}$ and render the images with ambient light and directional lighting.
@@ 3.1.2 Data Processing
3.2 Multi-view Image Generation and Pose Canonicalization
@@
3.2.1 IDUNet
Our IDUNet aims to retain sufficient features from the original posed image and ensure high consistency between the four generated views. Previous work, IP-Adapter [68], adds adapter modules into diffusion UNet structure. Appearance information in input images is transferred to the generated images via the cross-attention mechanism between the input image features and latent features. However, in practice, we observe that IP-Adapter cannot capture the fully detailed texture from input images, as such methods only utilize the global CLIP embeddings of the condition image, which loses pixel-level details during the image encoding, leading to inconsistent results.
To better incorporate features of the condition image, we propose IDUNet, to introduce pixel-level guidance in the generation process. Inspired by ControlNet [74], the structure of IDUNet is identical to the multi-view UNet. Unlike ControlNet, to ensure local patch-level interaction between all patches in both the denoised image and the condition image, we leverage the cross-attention between the latent tokens and condition image tokens rather than merely adding them together.
@@ 3.2.1 IDUNet
3.2.2 Multi-view UNet
The target of multi-view UNet is to generate multi-view A-pose images with highly consistent appearance from a single posed input image. Within the multi-view UNet, we simultaneously apply the denoising process on the four-view noisy latent $x_{\text{4v}}\in\mathrm{R}^{B\times 4\times N\times D}$. Here $B$, NN, and $D$ denote latent batch size, token numbers, and token feature dimension respectively. The transformer block in our multi-view UNet consists of a spatial self-attention module and a cross-attention module. As in MVDream [53], the spatial self-attention module directly takes tokens of all four noisy latent $x_{\text{4v}}$ and corresponding camera view embeddings. In the spatial self-attention layer, $x_{\text{4v}}$ is reshaped into $(B,4N,D)$ for patch-level cross-view interaction. This design allows the denoising UNet to capture the global relationships across different views, ensuring image generation with high consistency.
Figure 4: Pipeline for generating a final refined character mesh from generated multi-view images. In the first stage, we utilize a deep transformer-based network to generate a character with a coarse texture and then use a texture back-projection strategy to enhance the appearance of the generated mesh.
@@ 3.2.2 Multi-view UNet
$f_{\text{Cond}}$ $=\text{concat}(f_{\text{ID}},f_{\text{CLIP}})$ (1)
$x_{\text{4v}}$ $=\text{Cross\_Attn}(x_{\text{4v}},f_{\text{Cond}})$ (2)
Earlier work [27] on the diffusion process shows that applying zero sample-noise-ratio (SNR) at the last timestamp $T$ in training improves the final generation quality because the input noise in the inference stage is pure Gaussian noise. To achieve zero-SNR in the training stage, we manually set $SNR_{T}$ to zero and linearly scale other Guassian distribution parmaeters $\beta_{s}$. We set the UNet to directly output velocity $v_{\text{pred}}\in\mathrm{R}^{(B,4,N,D)}$ of the multi-view noisy latent and convert it to the noise $\epsilon_{\text{pred}}$. The final optimization target is given in Eq. 3.
@@ 3.2.3 Pose Canonicalization
Combined with IDUNet, our Multi-view diffusion model can successfully generate highly consistent orthographic view images while maintaining detailed features from the prompt image. To enable the diffusion model to achieve character pose canonicalization during the generation process, we jointly train the two UNets with the image pairs from our Anime3D dataset. However, simply training the diffusion network without extra pose constraints will lead to character layout misplacement and the emergence of unrelated body parts. To tackle these problems, we introduce character layouts to the diffusion models for generating A-pose character images by adopting OpenPose [5] to predict the pose embedding as an additional condition. The generated embeddings are directly added to the latent noise to aid the diffusion model in learning the relationships between character joints and the generated character layout. In the inference stage, we leverage three sets of OpenPose images from our Anime3D dataset and select the one with the highest CLIP score as the input pose condition.
3.3 3D Character Generation
@@
We now described how we efficiently generate 3D characters from the four-view images generated by our multi-view pose canonicalization diffusion model. As shown in Fig. 4, we adopt a coarse-to-fine process for the 3D character generation task. We first utilize a two-stage transformer-based network to reconstruct the geometry and coarse appearance of the character following the design of LRM [14]. Subsequently, we employ a texture back-projection strategy to quickly improve the texture quality using the generated high-resolution four-view images. Finally, we utilize Poisson Blending to reduce seams on the texture map.
3.3.1 Character Reconstruction with Coarse Texture

### v3

@@ Abstract.
Image-Driven Generation, 3D Avatar Generation, Avatar Pose Canonicalization
submissionid: 916copyright: rightsretainedjournal: TOGjournalvolume: 43journalnumber: 4article: 84journalyear: 2024publicationmonth: 7doi: 10.1145/3658217ccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Shape modelingccs: Computing methodologies Image manipulation
[…]
1. Introduction
The digital content industry’s rapid evolution has made the creation of high-quality 3D content a pivotal aspect across various domains, including film, video gaming, online streaming, and virtual reality (VR). Although manually modeled 3D content can attain exceptional quality, the significant time and labor investment required presents a substantial bottleneck. Addressing this, there has been a notable influx of exciting research (Tang et al., 2023a; Yoo et al., 2023; Raj et al., 2023; Qian et al., 2023; Wang and Shi, 2023; Deng et al., 2023; Melas-Kyriazi et al., 2023; Jun and Nichol, 2023; Nichol et al., 2022; Alwala et al., 2022; Wu et al., 2023a; Wang et al., 2021b; Wen et al., 2023; Liu et al., 2023d, b; Long et al., 2023) focused on generating 3D models from single images. This approach substantially lowers the barrier to entry for novice users, democratizing access to 3D content creation and potentially revolutionizing the field.
3D character models often feature complex articulations, leading to frequent self-occlusion in 2D images that pose significant challenges in reconstruction, generation, and animation. Moreover, these characters may assume a range of body poses, including some that are rare and challenging to accurately interpret, leading to a diverse yet imbalanced data domain. These complexities hinder the effective generation, rigging, and animating of such models. As a result, general 3D generation techniques (Poole et al., 2023; Chen et al., 2023a; Wang et al., 2023b; Lin et al., 2023a; Wang et al., 2023a; Metzer et al., 2023) and single-view 3D reconstruction methods (Thai et al., 2021; Hong et al., 2023b; Wang et al., 2021b; Wen et al., 2023; Wu et al., 2023a; Liu et al., 2023d; Alwala et al., 2022) often fall short in delivering optimal outcomes. Prior research (Cao et al., 2023; Huang et al., 2023a; Kolotouros et al., 2023; Huang et al., 2023b; Liao et al., 2023; Xiu et al., 2022, 2023; Saito et al., 2019, 2020; Zheng et al., 2022; Wang et al., 2023c) has explored the use of parametric models of 3D human bodies (Loper et al., 2015; Pavlakos et al., 2019; Alldieck et al., 2021; Li et al., 2017) as 3D priors. However, these methods are predominantly tailored to realistic human proportions and relatively tight clothing, limiting their applicability. This constraint is especially noticeable in the context of stylized characters, known for their exaggerated body proportions and complex clothing designs, which challenge the adaptability and effectiveness of these approaches.
@@ 1. Introduction
Our 3D character generation approach has two tightly interconnected stages: initially, lifting a single image to multiple viewpoints while simultaneously canonicalizing the input pose; following this, we proceed to reconstruct a 3D character using this canonical pose. This method is supported by two key insights: firstly, it incorporates established principles and successful techniques from recent advancements in controllable image generation (Zhang and Agrawala, 2023; Ye et al., 2023); secondly, it overcomes the challenges associated with sparse-view reconstruction for 3D characters. By focusing on the canonical pose, in which the geometric and texture structures are more clearly defined and self-occlusion is minimized, our approach simplifies the task of reconstructing both geometry and texture from limited views. The first stage involves a diffusion-based, image-conditioned multi-view generation model (Liu et al., 2023a; Wang and Shi, 2023; Shi et al., 2023a; Tang et al., 2023b), adept at capturing and translating both the global and local character features from the input image to the canonical pose, which further facilitates the generation of consistent canonical pose images across multiple views. The second stage employs a transformer-based, generalizable sparse-view reconstruction model (Hong et al., 2023b). This model is key to generating a coarsely textured 3D character model from the images produced in the first stage. We further refine the model’s texture resolution through projective texture mapping and Poisson Blending (Pérez et al., 2003), achieving a detailed final model. Furthermore, generating characters in a canonical pose also significantly benefits downstream applications, such as rigging and animation. We show the generated 3D characters with animation in Fig. 1. Our whole generation process takes less than 1 minute.
To train our pipeline, we have compiled a multi-pose, multi-view character dataset, focusing on anime characters due to their widespread availability online, notably on platforms such as VRoid Hub. We have amassed a collection of 13,746 characters and rendered these from various viewpoints across multiple body poses. This extensive collection has been organized into a dataset that we refer to as Anime3D.
@@ 2. Related Works
This section mainly discusses related works on diffusion-based 3D objects and avatar generation. Our CharacterGen also adopts transformer-based reconstruction models (Hong et al., 2023b; Li et al., 2023) for efficient 3D character generation. Space limitations preclude discussion of 3D human reconstruction works (Saito et al., 2019, 2020; Xiu et al., 2022, 2023; Cha et al., 2023).
2.1. Diffusion-Based 3D Object Generation
@@
Recently, diffusion methods have shown a strong ability to guide 3D object generation tasks in the past year. Pioneering work of DreamFusion (Poole et al., 2023) and SJC (Wang et al., 2023a) utilize score distillation sampling (SDS) to provide gradient guidance from pre-trained 2D diffusion models for text-to-3D generation tasks. Magic3D (Lin et al., 2023a) and Fantasia3D (Chen et al., 2023a) utilize an implicit tetrahedral field to support rendering with high resolution in the refinement stage. ProlificDreamer (Wang et al., 2023b) proposes VSD to distill gradient scores from a LoRA network to better learn the distribution of 3D objects. Zero123 (Liu et al., 2023c) presents a novel diffusion model to generate multi-view images that conform to the input image with given camera poses. Magic123 (Qian et al., 2023) combines both SDS and Zero123 guidance for generating 3D objects from image prompts, and adopts reconstruction loss to enhance front-view texture quality. MVDream (Shi et al., 2023b) and ImageDream (Wang and Shi, 2023) utilize multi-view diffusion models to provide highly consistent guidance in 3D object generation process. SyncDreamer (Liu et al., 2023a) utilizes 3D-aware attention modules to achieve synchronized multi-view image generation. Various other works (Jun and Nichol, 2023; Zhang et al., 2023b; Müller et al., 2023; Hui et al., 2022; Zeng et al., 2022; Nichol et al., 2022; Erkoç et al., 2023) employ 3D data to train diffusion models for direct 3D object generation, which are fast, but struggle with output diversity.
2.2. 3D Avatar Generation
Using strong human-body priors such as SMPL (Loper et al., 2015) and SMPL-X (Pavlakos et al., 2019), it is possible to generate high-quality human avatars based on the general 3D generation methods. EVA3D (Hong et al., 2023a) combines a GAN backbone with a pose-guided sampling method to generate high-quality 3D human avatars. AvatarCLIP (Hong et al., 2022) first solves text-to-human generation by utilizing the pre-trained CLIP model to guide optimization of geometry and color networks.
@@
Dreamavatar (Cao et al., 2023) and AvatarCraft (Jiang et al., 2023) utilize SMPL to initialize the implicit human geometry used in the diffusion-guided generation process. DreamHuman (Kolotouros et al., 2023) adopts ImGHum (Alldieck et al., 2021) as body priors and proposes a focus rendering mechanism to better reconstruct the detailed geometry of avatars. DreamWaltz (Huang et al., 2023a) utilizes ControlNet (Zhang and Agrawala, 2023) to provide pose guidance to finetuning the animated representation. AvatarVerse (Zhang et al., 2023a) and AvatarStudio (Zhang et al., 2023c) both utilize a DensePose-guided ControlNet in the generation process to circumvent the multi-face “Janus” problem and to support part geometry optimization. TeCH (Huang et al., 2023b) supports image-prompt avatar generation by training an additional DreamBooth model (Ruiz et al., 2023) on the input image for the SDS guidance model. TADA (Liao et al., 2023) directly distills 2D diffusion models to optimize the normals and displacements of an SMPL body mesh.
@@ 3.1. Anime3D Dataset
3.1.1. Data Acquisition
Existing large 3D object datasets like Objaverse (Deitke et al., 2023) or OmniObject3D (Wu et al., 2023b) do not contain enough 3D stylized characters for our training purposes. Inspired by PAniC3D (Chen et al., 2023b), we first collected a large dataset of nearly 14,500 anime characters from the VRoidHub (VRoid, 2022) and then removed non-humanoid data, to leave 13,746 character models.
Figure 2. An example character from our Anime3D dataset from four different camera views, demonstrates how we organize the image pairs during training to extend UNet’s ability to determine a canonical pose.
Figure 3. Pipeline for generating four views of consistent images, showing how our IDUNet extracts local pixel-level features to strengthen the multi-view UNet. Here ”Q”, ”K”, and ”V” denote the query, key, and value matrix in the attention mechanism.
3.1.2. Data Processing
We need to render all the objects into image format to fine-tune 2D diffusion models. We utilize the three-vrm (Pixiv, 2019) framework to render these characters.
We first obtain “A-pose” characters and posed characters to generate canonical pose and random pose image pairs. For A-pose characters, we set the joint rotation of the left and right arm to $45^{\circ}$ in the $Z$-axis and the left and right upper leg to $6^{\circ}$ in the $Z$-axis. All other joint parameters remain untouched. For the posed character setting, we download 10 human skeleton animations from Mixamo (MixamoInc., 2009), including sitting, singing, and walking, etc. We randomly select frames from these animations and apply the corresponding motion to the VRM character models. In addition, we also randomize joints of the mouth and eyes to generate a variety of facial expressions, such as winking. We normalize the bounding boxes of the character model to $[-0.5,0.5]^{3}$. We configure the camera field of view (FoV) to $40^{\circ}$ and set the distance between the camera and the scene origin to 1.5. The character images are rendered with ambient light and directional lighting.
@@ 3.1.2. Data Processing
3.2. Multi-view Image Generation and Pose Canonicalization
@@
3.2.1. IDUNet
Our IDUNet aims to retain sufficient features from the original posed image and ensure high consistency between the four generated views. Previous work, IP-Adapter (Ye et al., 2023), adds adapter modules into diffusion UNet structure. Appearance information in input images is transferred to the generated images via the cross-attention mechanism between the input image features and latent features. However, in practice, we observe that IP-Adapter cannot capture the fully detailed texture from input images, as such methods only utilize the global CLIP embeddings of the condition image, which loses pixel-level details during the image encoding, leading to inconsistent results.
To better incorporate features of the condition image, we propose IDUNet, to introduce pixel-level guidance in the generation process. Inspired by ControlNet (Zhang and Agrawala, 2023), the structure of IDUNet is identical to the multi-view UNet. IDUNet takes a fixed text-prompt ”best quality” to provide general guidance in the generation process. Unlike ControlNet, to ensure local patch-level interaction between all patches in both the denoised image and the condition image, we leverage the cross-attention between the latent tokens and condition image tokens rather than merely adding them together.
@@ 3.2.1. IDUNet
3.2.2. Multi-view UNet
The target of multi-view UNet is to generate multi-view A-pose images with highly consistent appearance from a single posed input image. Within the multi-view UNet, we simultaneously apply the denoising process on the four-view noisy latent $x_{\text{4v}}\in\mathrm{R}^{B\times 4\times N\times D}$. Here $B$, $N$, and $D$ denote latent batch size, token numbers, and token feature dimension respectively. The multi-view UNet takes the extrinsic camera matrices of the four views as spatial guidance. During the inference stage, the camera poses are set with the fixed azimuths of $\{0^{\circ},90^{\circ},180^{\circ},270^{\circ}\}$ and the elevation of $0^{\circ}$. The transformer block in our multi-view UNet consists of a spatial self-attention module and a cross-attention module. As in MVDream (Shi et al., 2023b), the spatial self-attention module directly takes tokens of all four noisy latent $x_{\text{4v}}$ and corresponding camera view embeddings. In the spatial self-attention layer, $x_{\text{4v}}$ is reshaped into $(B,4N,D)$ for patch-level cross-view interaction. This design allows the denoising UNet to capture the global relationships across different views, ensuring image generation with high consistency.
Figure 4. Pipeline for generating a final refined character mesh from generated multi-view images. In the first stage, we utilize a deep transformer-based network to generate a character with a coarse texture and then use a texture back-projection strategy to enhance the appearance of the generated mesh.
@@ 3.2.2. Multi-view UNet
(1) $f_{\text{Cond}}$ $=\text{concat}(f_{\text{ID}},f_{\text{CLIP}})$
@@
Earlier work (Lin et al., 2023b) on the diffusion process shows that applying zero sample-noise-ratio (SNR) at the last timestamp $T$ in training improves the final generation quality because the input noise in the inference stage is pure Gaussian noise. To achieve zero-SNR in the training stage, we manually set $SNR_{T}$ to zero and linearly scale other Guassian distribution parmaeters $\beta_{s}$. We set the UNet to directly output velocity $v_{\text{pred}}\in\mathrm{R}^{(B,4,N,D)}$ of the multi-view noisy latent and convert it to the noise $\epsilon_{\text{pred}}$. The final optimization target is given in Eq. 3.
(3) $L_{\text{4v}}=||\epsilon_{\text{4v}}-\epsilon_{\text{pred}}||^{2}_{2}.$
@@ 3.2.3. Pose Canonicalization
Combined with IDUNet, our Multi-view diffusion model can successfully generate highly consistent orthographic view images while maintaining detailed features from the prompt image. To enable the diffusion model to achieve character pose canonicalization during the generation process, we jointly train the two UNets with the image pairs from our Anime3D dataset. However, simply training the diffusion network without extra pose constraints will lead to character layout misplacement and the emergence of unrelated body parts. To tackle these problems, we introduce character layouts to the diffusion models for generating A-pose character images by adopting OpenPose (Cao et al., 2021) to predict the pose embedding as an additional condition. The generated embeddings are directly added to the latent noise to aid the diffusion model in learning the relationships between character joints and the generated character layout. Since characters have various body shapes, we leverage three different sets of OpenPose images from our Anime3D dataset and select the one with the highest CLIP score as the input pose condition in the inference stage.
3.3. 3D Character Generation
@@
We now describe how we efficiently generate 3D characters from the four-view images generated by our multi-view pose canonicalization diffusion model. As shown in Fig. 4, we adopt a coarse-to-fine process for the 3D character generation task. We first utilize a two-stage transformer-based network to reconstruct the geometry and coarse appearance of the character following the design of LRM (Hong et al., 2023b). Subsequently, we employ a texture back-projection strategy to quickly improve the texture quality using the generated high-resolution four-view images. Finally, we utilize Poisson Blending (Pérez et al., 2003) to reduce seams on the texture map.
3.3.1. Character Reconstruction with Coarse Texture